# Bitune: Bidirectional Instruction-Tuning

### v1

@@ Abstract
We introduce Bitune, a method that improves instruction-tuning of pretrained decoder-only large language models, leading to consistent gains on downstream tasks. Bitune applies both causal and bidirectional attention to the prompt, to obtain a better representation of the query or instruction. We realize this by introducing two sets of parameters, for which we apply parameter-efficient finetuning techniques. These causal and bidirectional features are then combined into a weighted average with trainable coefficients, which is subsequently used to generate new tokens. We demonstrate significant improvements in zero-shot performance on commonsense reasoning, arithmetic, and language understanding tasks, while extensive ablation studies validate the role of each component and demonstrate the method’s agnosticism to different PEFT techniques.
@@ 1 Introduction
Large Language models (LLMs) are being used in many practical scenarios where humans interact with them through questions and answers. In applications such as general purpose assistants (OpenAI, 2024a), medical diagnosticians (Thirunavukarasu et al., 2023), game-conversation generation (Cox & Ooi, 2023) or coding-assistants (Roziere et al., 2023), the ability for an LLM to precisely follow an instruction and answer a question are of primary concern.
@@
Correspondingly, Instruction-Tuning (IT)  (Chung et al., 2024; Ouyang et al., 2022a) is the prevailing paradigm for finetuning LLMs after their self-supervised pretraining phase to improve them for such tasks. Here, the model is trained on a dataset comprised of pairs of instructions and corresponding responses. Given the instruction-with-response structure of IT data, the generation of an LLM response can be divided into two phases: first, converting the instruction into Key and Value (KV) embeddings, stored in a KV-cache, which we refer to as instruction features; second, using these features to autoregressively generate an answer. Due to this task’s inherently conditional nature, the instruction features’ effectiveness is crucial for obtaining high-quality model outputs.
@@
In the past, bidirectional attention (Schuster & Paliwal, 1997) has been a key technique for obtaining stronger features for words or tokens. This is because the meaning of a word depends greatly on its context. In particular, for some words in a sentence, the information that comes later might be far more informative for generating a meaningful representation and resolving ambiguities. But with only uni-directional causal attention, where the representation of each word is restricted to depend solely on the words that came before, this cannot be achieved. This is the reason why many previous transformers such as encoder-only BERT (Devlin et al., 2019) and encoder-decoder T5 (Raffel et al., 2020) employed bidirectional attention to improve the encoding of the input and why tasks like text retrieval (Lewis et al., 2020; Li & Li, 2023) and even the latest text-to-image generative models (OpenAI, 2024b; Esser et al., 2024) still rely on this.
@@
However, in the context of LLMs, architectures utilizing bidirectional attention have fallen out of favor, as decoder-only models such as GPT (OpenAI, 2024a) and Llama (AI@Meta, 2024) have focused on and vastly improved the generative performance of language models. These architectures are trained by large volumes of data with next-token prediction, eschewing any look-ahead mechanism for the sake of better autoregressive modeling. As there is simply more unlabeled data available for pretraining, training a decoder-only architecture on unlabeled data, and then finetuning it for tasks with instruction-tuning, is the best modus operandi of today (Wang et al., 2022). However, with this switch to decoder-only architectures, we lost bidirectional attention in the process. As we know this can improve feature representations for instructions, we set out to re-introduce bidirectional attention, such that it can be integrated into pretrained decoder-only LLMs.
@@
  • We propose a novel method, Bitune, that improves the performance of pretrained decoder-only LLMs in question-answering scenarios.
@@
  • We conduct an extensive ablation study investigating the necessity of each component of the method, and showing the method’s PEFT-agnosticism.
@@ 2 Bidirectional Instruction-Tuning
Algorithm 1 Python-like pseudocode of Bitune inference.
@@
# theta - tensor with mixing coeff. for each layer
# theta_init - initial value of mixing coefficients
@@
c_token = SEP # Initialize generation with a predefined token
@@
while c_token != EOS: # Stop generation on the end-of-sequence token
# Get features of current token and logits of next token
@@
kv = concat(kv, (k, v)) # Concatenate obtained features with current KV cache
c_token = get_token(logits) # Determine next predicted token based on logits
answer.append(c_token) # Append generated token to the answer
@@
where $|a_{i}|$ denotes the length of the answer. Note how compared to the regular language modeling objective, the response is already conditional (on the instruction $q$) even for the first generated token.
@@
In this work, we introduce, Bitune, a method to leverage this two-phase process to improve instruction-tuning of language models. In our approach, the model processes the instruction with both causal and bidirectional attention using separate sets of parameters, leading to an enhanced KV-cache that is then used to condition the answer. Figure 1 provides an overview of the method, while Algorithm 1 presents pseudocode for the inference process.
@@
$\text{Attention}(Q,K,V,M)=\text{softmax}\left(\frac{QK^{T}}{\sqrt{d_{k}}}+M\right)V,$ (3)
where $Q$ are the queries, $M$ is the attention mask, and $d_{k}$ is the dimension of keys and queries.
For the causal pass, the mask $M_{\color[rgb]{1,0,0}c}$ enforces causality by masking future tokens, such that tokens $j$ can only attend to earlier tokens $i\leq j$, while for the bidirectional pass, no masking is applied:
$M_{{{\color[rgb]{1,0,0}c}}}(i,j)=\begin{cases}0&\text{if }i\leq j\\ -\infty&\text{if }i>j\end{cases},\qquad M_{\color[rgb]{0,0,1}b}(i,j)=0.$ (4)
@@
$K_{\texttt{Bitune}}=K_{\color[rgb]{1,0,0}c}\cdot(1-\alpha_{k})+K_{\color[rgb]{0,0,1}b}\cdot\alpha_{k},\quad V_{\texttt{Bitune}}=V_{\color[rgb]{1,0,0}c}\cdot(1-\alpha_{v})+V_{\color[rgb]{0,0,1}b}\cdot\alpha_{v},$ (5)
@@
$\alpha_{j}=\frac{|\theta_{j}|}{\theta_{\text{init}}+|\theta_{j}|},\quad j\in\{k,v\}$ (6)
@@
Instead, we adapt our model to use bidirectional attention by using parameter-efficient finetuning methods. These introduce only a fraction of trainable parameters, making it viable to have two modified variants of the model within a single forward pass. In the default configuration of our method, we utilize the Low-Rank Adaptation (LoRA) of Hu et al. (2022) to adapt the model. However, our method is PEFT-agnostic and can also utilize different methods for updating the weights, as demonstrated in our ablations section.
In summary, Bitune is a method for improved instruction-tuning that learns a combination of causal and bidirectional features by applying two sets of lightweight PEFT methods to a pretrained model.
@@ 3 Experiments
PIQA ARC CSQA SIQA MMLU Avg.
Model Method
Gemma-2B Pretrained 57.5 36.9 35.5 38.2 34.0 40.4
@@
3.1 Instruction-Tuning
Specifically, we use a subset of the cleaned UltraFeedback (Cui et al., 2023) dataset, which contains instructions and corresponding answers generated by various LLMs. From this dataset, we select completions generated by GPT-4 (OpenAI, 2024a), ensuring high-quality responses for training. To fit every model on a single GPU, we filter out samples longer than 512 tokens, which leaves us with around 10,000 samples for training.
We test the method on pretrained decoder-only language models of two different scales of approximately 2 billion and 7 billion parameters. The specific models used in our experiments are: Gemma 2B and 7B (Gemma Team et al., 2024), Llama2 7B (Touvron et al., 2023) and Llama3 8B (AI@Meta, 2024), and Phi-2 (Li et al., 2023), which has 2.7 billion parameters. We use HuggingFace Transformers (Wolf et al., 2020) implementation of these models. We limit our experiments to these specific models due to their popularity and also our limited resources.
[…]
For each model, we tune the learning rate on the LoRA baseline using steps on the approximate logarithmic scale ($1\mathrm{e}{-4}$, $3\mathrm{e}{-4}$, $1\mathrm{e}{-3}$, $3\mathrm{e}{-3}$), and then apply the same rate to the other approaches. Note that this potentially puts our method at a disadvantage compared to the LoRA baseline. All hyperparameters are reported in the Appendix 6.2.
@@ 3.1 Instruction-Tuning
In the decoding phase of the inference with Bitune, to initiate generation, the model requires at least a single token to obtain a set of attention queries, in addition to the keys and values extracted from the instruction. To facilitate this, one can introduce a new learnable <sep> token that would be placed at the beginning of modeled answer, or utilize an existing token. For our experiments, we opted to move the last token of the instruction template to the beginning of the modeled answer. For details on the instruction template used, please refer to the section 6.5 of the Appendix.
@@ 3.2 Downstream Task Training
This complementary experiment verifies if Bitune increases the capacity of the model within the narrow scope of a single task. It follows the setup from the instruction-tuning experiments with a few changes. Namely, models are not instruction-tuned but trained separately for each evaluation task using the corresponding training set. We use PIQA, ARC, CSQA, and SIQA introduced earlier, and an additional arithmetic task, GSM8K (Cobbe et al., 2021).
@@ Results.
Table 2: Result for the downstream task training. We show accuracy on downstream tasks for the baseline LoRA finetuning and Bitune, averaged over 3 seeds. On average, our method works better than standard LoRA. We see the most significant gains on the GSM8K dataset, but slightly lower results for Gemma-7B and the SIQA task.
PIQA ARC CSQA SIQA GSM8K Avg.
Model Method
Gemma-2B LoRA 81.4 58.0 77.2 77.4 30.2 64.8
@@
Table 2 presents the results, demonstrating improvements when finetuning on the downstream tasks with Bitune, similar to those seen with instruction-tuning. While there are a few cases where the baseline finetuning achieves better results on specific tasks, when considering the average gains, applying our method is beneficial across all models. Most importantly, on the GSM8K dataset, we see consistently high gains, suggesting that our method improves the model’s reasoning ability in generative tasks. The qualitative example for instruction-tuned model in Table 6 also supports this (more examples in Appendix 6.7). Similar to the instruction-tuning results, the highest gains are observed on the Gemma-2B model, while the lowest on the Gemma-7B. This indicates that the effectiveness of our method depends on the specific model used.
@@ 3.3 Ablations
Table 3: Ablation study on components of Bitune. We report zero-shot accuracy averaged over PIQA, ARC, CSQA, SIQA and MMLU tasks. The components are explained in section 3.3. We see that all ablated variants outperform the LoRA baseline, and combining all components performs the best. Of note, especially bidirectional attention improves results the most.
Avg. Acc.
Model Method Causal Bidir. Mixing Sep. Weights
Gemma-2B LoRA - - - 45.7
@@
The results, averaged over three seeds and presented in Table 3, indicate that all variants of the method lead to gains over the baseline LoRA finetuning. However, the highest gains are observed in the full variant of Bitune, demonstrating that each component contributes to the method’s effectiveness.
@@
PIQA ARC CSQA SIQA MMLU Avg.
Model Method
Gemma-2B LoRA 66.7 43.4 42.3 44.3 31.7 45.7
@@
The results are shown in Table 4. We find consistent gains across all three PEFT methods we analyze, with gains ranging from $+1.6\%$ to $+4.0\%$ for averaged accuracies. This demonstrates that Bitune is PEFT-agnostic and can be combined with existing and future innovations in PEFT methods.
Initialization of Mixing Coefficient

### v2

# Bitune: Leveraging Bidirectional Attention to Improve Decoder-Only LLMs

@@ Abstract
Decoder-only large language models typically rely solely on masked causal attention, which limits their expressiveness by restricting information flow to one direction. We propose Bitune, a method that enhances pretrained decoder-only LLMs by incorporating bidirectional attention into prompt processing. We evaluate Bitune in instruction-tuning and question-answering settings, showing significant improvements in performance on commonsense reasoning, arithmetic, and language understanding tasks. Furthermore, extensive ablation studies validate the role of each component of the method, and demonstrate that Bitune is compatible with various parameter-efficient finetuning techniques and full model finetuning.
Bitune: Leveraging Bidirectional Attention to Improve Decoder-Only LLMs
Dawid J. Kopiczkothanks: Correspondence: dawid.kopiczko@utn.de Fundamental AI Lab University of Technology Nuremberg          Tijmen Blankevoort Meta          Yuki M. Asano Fundamental AI Lab University of Technology Nuremberg
@@ 1 Introduction
Large Language Models (LLMs) are being deployed in numerous practical applications where humans engage with them through various forms of natural language interaction. In use cases such as general purpose assistants (OpenAI, 2024), medical diagnosticians (Thirunavukarasu et al., 2023), game-conversation generation (Cox and Ooi, 2023) or coding-assistants (Roziere et al., 2023), the ability for an LLM to precisely interpret and respond to user inputs is of primary concern.
@@
Correspondingly, Instruction-Tuning (IT)  (Chung et al., 2024; Ouyang et al., 2022a) is the prevailing paradigm for finetuning LLMs after their self-supervised pretraining phase to improve them for such tasks. Here, the model is trained on a dataset comprised of pairs of instructions and corresponding responses. Given the instruction-with-response structure of IT data, the generation of an LLM response can be divided into two phases: first, converting the instruction into key and value embeddings, which we refer to as instruction features; second, using these features to autoregressively generate an answer. Due to this task’s inherently conditional nature, the instruction features’ effectiveness is crucial for obtaining high-quality model outputs.
@@
In the past, bidirectional attention (Schuster and Paliwal, 1997) has been a key technique for obtaining stronger features for words or tokens. This is because the meaning of a word depends greatly on its context. In particular, for some words in a sentence, the information that comes later might be far more informative for generating a meaningful representation and resolving ambiguities. With only uni-directional causal attention, where the representation of each word is restricted to depend solely on the words that came before, this cannot be achieved. This is the reason why many previous transformers such as encoder-only BERT (Devlin et al., 2019) and encoder-decoder T5 (Raffel et al., 2020) employed bidirectional attention to improve the encoding of the input and why tasks like text retrieval (Lewis et al., 2020; Li and Li, 2023) still rely on this.
@@
However, in the context of LLMs, architectures utilizing bidirectional attention have fallen out of favor, as decoder-only models such as GPT (OpenAI, 2024) and Llama (AI@Meta, 2024) have focused on and vastly improved the generative performance of language models. These architectures are trained by large volumes of data with next-token prediction, eschewing any look-ahead mechanism for the sake of better autoregressive modeling. As there is simply more unlabeled data available for pretraining, training a decoder-only architecture on unlabeled data, and then finetuning it for tasks with instruction-tuning, is the best modus operandi of today (Wang et al., 2022). However, with this switch to decoder-only architectures, we lost bidirectional attention in the process. As we know this can improve feature representations for instructions, we set out to re-introduce bidirectional attention, such that it can be integrated into pretrained decoder-only LLMs.
@@
  • We propose a novel method, Bitune, that improves the performance of pretrained decoder-only LLMs in instruction-following and question-answering settings.
@@
  • We conduct an extensive ablation study investigating the necessity of each component of the method, and showing the method’s PEFT-agnosticism, as well as its effectiveness in full finetuning scenarios.
@@ 2 Bidirectional Instruction-Tuning
Algorithm 1 Pseudocode of Bitune inference.
@@
# theta - tensor with mixing coefficients
# theta_init - initial value of theta
# Extract causal features
k_c, v_c = model_causal(prompt)
# Extract bidirectional features
k_b, v_b = model_bidir(prompt)
@@
# Initialize generation with a predefined token
c_token = SEP
@@
# Stop generation at the end-of-sequence token
while c_token != EOS:
# Get features of current token and logits
# of next token
@@
# Concatenate it with current KV cache
kv = concat(kv, (k, v))
# Determine next predicted token
c_token = get_token(logits)
# Append generated token to the answer
answer.append(c_token)
@@
where $|a|$ denotes the length of the answer. Note how compared to the regular language modeling objective, the response is already conditional (on the instruction $q$) even for the first generated token.
@@
In this work, we introduce Bitune, a method to leverage this two-phase process to improve instruction-tuning of language models. In our approach, the model processes the instruction with both causal and bidirectional attention using separate sets of parameters, leading to an enhanced KV-cache that is then used to condition the answer. Figure 1 provides an overview of the method, while Algorithm 1 presents pseudocode for the inference process.
@@
$\text{Attn}(Q,K,V,M)=\sigma({QK^{T}}/{\sqrt{d_{k}}}+M)V,$ (3)
where $\sigma$ is the softmax, $Q$ are the queries, $M$ is the attention mask, and $d_{k}$ is the dimension of keys and queries. For the causal pass, the mask $M_{\color[rgb]{1,0,0}c}$ enforces causality by masking future tokens, such that tokens $j$ can only attend to earlier tokens $i\leq j$, while for the bidirectional pass, no masking is applied:
$M_{{{\color[rgb]{1,0,0}c}}}(i,j)$ $=\begin{cases}0&\text{if }i\leq j\\ -\infty&\text{if }i>j\end{cases}$ (4)
$M_{\color[rgb]{0,0,1}b}(i,j)$ $=0$ (5)
@@
$K_{\texttt{Bitune}}$ $=K_{\color[rgb]{1,0,0}c}\cdot(1-\alpha_{k})+K_{\color[rgb]{0,0,1}b}\cdot\alpha_{k}$ (6)
$V_{\texttt{Bitune}}$ $=V_{\color[rgb]{1,0,0}c}\cdot(1-\alpha_{v})+V_{\color[rgb]{0,0,1}b}\cdot\alpha_{v},$ (7)
@@
$\alpha_{j}={|\theta_{j}|}/({\theta_{\text{init}}+|\theta_{j}|}),\quad j\in\{k,v\}$ (8)
@@
Instead, we adapt our model using parameter-efficient finetuning methods. These introduce only a fraction of trainable parameters, making it viable to have two modified variants of the model within a single forward pass. In the default configuration of our method, we utilize the Low-Rank Adaptation (LoRA) of Hu et al. (2022) to adapt the model. However, Bitune can utilize different methods for updating the weights, including full model finetuning and other parameter-efficient techniques, as demonstrated in our ablations section.
@@ 3 Experiments
3.1 Instruction-Tuning
Specifically, we use a subset of the cleaned UltraFeedback (Cui et al., 2023) dataset, which contains instructions and corresponding answers generated by various LLMs. From this dataset, we select completions generated by GPT-4 (OpenAI, 2024), ensuring high-quality responses for training. To fit every model on a single GPU, we filter out samples longer than 512 tokens, which leaves us with roughly 10,000 samples for training. For results on another instruction-tuning dataset, please see the Appendix A.6.
We test the method on pretrained decoder-only language models of two different scales of approximately 2 billion and 7 billion parameters. The specific models used in our experiments are: Gemma 2B and 7B (Gemma Team et al., 2024), Llama2 7B (Touvron et al., 2023) and Llama3 8B (AI@Meta, 2024), and Phi-2 (Li et al., 2023), which has 2.7 billion parameters. We use HuggingFace Transformers (Wolf et al., 2020) implementation of these models.
[…]
For each model, we tune the learning rate on the LoRA baseline using steps on the approximate logarithmic scale ($1\mathrm{e}{-4}$, $3\mathrm{e}{-4}$, $1\mathrm{e}{-3}$, $3\mathrm{e}{-3}$), and then apply the same rate to the other approaches. Note that this potentially puts our method at a disadvantage compared to the LoRA baseline. All hyperparameters are reported in the Appendix A.2.
@@ 3.1 Instruction-Tuning
Method PIQA ARC CSQA SIQA MMLU Avg.
Gemma-2B
Pretrained 57.5 36.9 35.5 38.2 34.0 40.4
@@
In the decoding phase of the inference with Bitune, to initiate generation, the model requires at least a single token to obtain a set of attention queries, in addition to the keys and values extracted from the instruction. To facilitate this, one can introduce a new learnable <sep> token that would be placed at the beginning of modeled answer, or utilize an existing token. For our experiments, we opted to move the last token of the instruction template to the beginning of the modeled answer. For details on the instruction template used, please refer to the section A.10 of the Appendix.
@@ 3.2 Downstream Task Training
This complementary experiment verifies whether Bitune increases the capacity of the model within the narrow scope of a single task. It follows the setup from the instruction-tuning experiments with a few changes. Namely, models are not instruction-tuned but trained separately for each evaluation task using the corresponding training set. We use PIQA, ARC, CSQA, and SIQA introduced earlier, and an additional arithmetic task, GSM8K (Cobbe et al., 2021).
@@ Results.
Table 2: Results for the downstream task training. We show accuracy on downstream tasks for the baseline LoRA finetuning and Bitune, averaged over 3 seeds. On average, our method works better than standard LoRA. We see the most significant gains on the GSM8K dataset, but slightly lower results for Gemma-7B and the SIQA task.
Method PIQA ARC CSQA SIQA GSM8K Avg.
Gemma-2B
LoRA 81.4 58.0 77.2 77.4 30.2 64.8
[…]
@@
Table 2 presents the results, demonstrating improvements when finetuning on the downstream tasks with Bitune, similar to those seen with instruction-tuning. While there are a few cases where the baseline finetuning achieves better results on specific tasks, when considering the average gains, applying our method is beneficial across all models. Most importantly, on the GSM8K dataset, we see consistent high gains, suggesting that our method improves the model’s reasoning ability in generative tasks. We present additional results on GSM8K with a 22B parameter model in the Appendix A.7.
Similar to the instruction-tuning results, the highest gains are observed on the Gemma-2B model, while the lowest on the Gemma-7B. This indicates that the effectiveness of our method depends on the specific model used.
3.3 Chain-of-Thought Reasoning
To further test the effectiveness of our method on generative tasks that require explicit reasoning, we created four additional training sets for PIQA, ARC, CSQA, and GSM8K. Each set was distilled from the GPT-4.1 model (OpenAI, 2025) by prompting it to solve every training-split question step-by-step; the resulting reasoning traces were then used to finetune Gemma-2B and Llama-3-8B on each benchmark separately. At evaluation time the models generated a full answer, including intermediate reasoning, and scores were averaged over three random seeds.
The original versions of these benchmarks supply only the final label, forcing a model trained on them to perform all reasoning in latent space and yielding short generations that do not reflect real-world usage. Our distilled chain-of-thought data instead mirrors the way large language models are typically invoked, producing longer, multi-step explanations before the answer, thereby providing a more faithful test bed. Moreover, because the public test sets may already appear in pre-training corpora, training the models to answer the same questions in a custom, step-by-step format reduces the risk that apparent gains stem from data contamination. The prompts used to distill the datasets, along with example completions, are available in Appendix A.9.
Results.
Table 3 shows that Bitune improves performance on chain-of-thought reasoning tasks, raising average accuracy by 2.4 pp for Gemma-2B and 1.3 pp for Llama-3-8B.
Table 3: Results for the chain-of-thought training. Models were trained on downstream tasks with chain-of-thought reasoning traces distilled from GPT-4.1 model, and evaluated by generating the reasoning steps before providing the final answer to given question. Across the tasks and models tested, Bitune leads to significant gains when compared to LoRA baselines.
Method PIQA ARC CSQA GSM8K Avg.
Gemma-2B
LoRA 77.5 55.4 63.2 40.3 59.1
[…]
Bitune 78.4 56.6 67.1 43.7 61.5
Llama3-8B
LoRA 88.2 82.5 78.6 70.3 79.9
[…]
Bitune 89.6 82.8 79.0 73.5 81.2
3.4 Inference Speed Comparison
During inference, Bitune performs two forward passes over the instruction sequence and merges the resulting features, which affects prefilling runtime. We compare inference time for Gemma-2B on a long-context instruction of 2000 tokens and report the time required to prefill this instruction and to generate 2000 subsequent tokens on a single A100 GPU.
Table 4: Inference-time comparison for Gemma-2B. We report the time required to prefill a 2000-token instruction and to generate a further 2000 tokens on a single A100 GPU.
Method Prefill Generation
LoRA 0.03s 62.15s
Bitune 0.28s 62.28s
Table 4 shows that the extra latency for processing the instruction is negligible because most of the computation time is spent on autoregressive answer generation.
@@ 3.5 Ablations
Table 5: Ablation study on components of Bitune. We report zero-shot accuracy averaged over PIQA, ARC, CSQA, SIQA and MMLU tasks. The components are explained in section 3.5. We see that all ablated variants outperform the LoRA baseline, and combining all components performs the best.
Method Causal Bidir. Mixing Sep. weights Acc.
Gemma-2B
LoRA - - - 45.7
@@
The results, averaged over three seeds and presented in Table 5, indicate that all variants of Bitune lead to gains over the baseline LoRA finetuning. However, the highest gains are observed in the full variant of Bitune, demonstrating that each component contributes to the method’s effectiveness.
@@
Method PIQA ARC CSQA SIQA MMLU Avg.
Gemma-2B
LoRA 66.7 43.4 42.3 44.3 31.7 45.7
@@
The results are shown in Table 6. We find consistent gains across all three PEFT methods we analyze, with gains ranging from $1.6\%$ to $4.0\%$ for averaged accuracy. This demonstrates that Bitune is PEFT-agnostic and can be combined with existing and future innovations in PEFT methods.
Full Finetuning
Additionally, we test whether Bitune leads to gains with full finetuning (Full-FT), by optimizing two sets of full model’s parameters. We conduct experiments on Gemma-2B model, and compare results with standard Full-FT baseline. The results in Table 7 demonstrate that Bitune improves the model’s performance even in full finetuning scenarios.
Table 7: Results for full finetuning Gemma-2B model on the instruction-tuning setup. We compare our method to full finetuning baseline. For Bitune, we optimize two sets of full model’s weights.
Method PIQA ARC CSQA SIQA MMLU Avg.
Full FT 69.0 46.7 43.5 43.7 34.8 47.5
[…]
Attention Mask of Second Pass
Table 8: Ablation of Bitune’s attention. We vary the attention mask for extracting the second set of features.
Model Attention Mask Avg. Acc
Gemma-2B Causal 46.9
Anti-causal 46.7
Bidirectional 49.7
Llama3-8B Causal 71.1
Anti-causal 71.4
Bidirectional 72.7
Results shown in Table 8 indicate that the instruction has to be processed with full bidirectional attention to achieve the highest gains. Combining causal and anti-causal features independently does not lead to the same high performance.
4 Related Work
Additionally, UL2 (Tay et al., 2023) introduces a pretraining objective called "Mixture of Denoisers", which combines various denoising strategies, including the prefix-LM approach. Lastly, XLNet (Yang et al., 2019) also allows for non-causal word ordering by allowing random permutations to be used with a next-token prediction objective.
Bitune vs. Prefix-LM.
Both Bitune and the classic prefix-LM masking scheme grant full bidirectional access to the instruction prefix. The key difference is that Bitune is a hybrid approach designed for the optimal adaptation of decoder-only models. Existing large language models are almost always pre-trained as pure decoder-only networks; simply replacing the causal mask with the prefix-LM masking scheme before finetuning therefore discards the causal features encoded during pretraining and forces the model to relearn the mechanics of bidirectional attention over the prefix tokens. Bitune keeps the original causal stream intact and adds a second bidirectional pass, injecting fresh capacity into the model. Finetuning then proceeds in a single architecture that unifies the two attention regimes. Our ablation study (Section 3.5) confirms the payoff: Bitune yields larger gains than straightforward “prefix-LM conversion” (ablation Naive Bidir). Whether the same advantage holds when both schemes are trained from scratch remains an open question.
5 Conclusion
This work proposes a method that exploits the inherent instruction–answer structure of IT datasets to incorporate bidirectional attention into pretrained decoder-only models. Bitune demonstrates general applicability across different models and scales, and it consistently improves performance across commonsense reasoning, arithmetic, and language understanding tasks. We further demonstrate that the method is compatible with different existing PEFT methods and will likely benefit from further developments in this area.
Limitations
During standard instruction-tuning training, the instruction and the answer are processed in a single forward pass. In our method, this processing is explicitly split into phases - extracting instruction features with causal attention, bidirectional attention, and answer modeling, which results in increased both training time and memory usage.
However, this is a minor limitation in the context of instruction-tuning, since typically smaller datasets are used compared to pretraining, leading to relatively short training times. Furthermore, as shown in Section 3.4, the additional latency during prefilling at inference time is negligible, since most time is spent on sequential token generation.
Appendix A Appendix
A.1 Pseudocode for Bitune Training Step
Algorithm 2 Python-like pseudocode of Bitune Training Step.
# prompt - tensor with tokenized instruction
# answer - tensor with tokenizer answer to model
# theta - trainable tensor with mixing coeff. for each layer
# theta_init - initial value of mixing coefficients
# Combine both sets of features
alpha = theta.abs() / (theta.abs() + theta_init)
k = k_c * (1 - alpha) + k_b * alpha
v = v_c * (1 - alpha) + v_b * alpha
kv = (k, v)
logits = model_causal(answer, kv)
loss = compute_loss(logits, answer)
loss.backward()
update_parameters(model_causal, model_bidir, theta)
A.2 Hyperparameters
Table 9: Hyperparameters shared across models and datasets.
Hyperparameter Value
GPUs 1
Optimizer AdamW
LR Scheduler Linear
Weight Decay 0.0
Batch Size (incl. accumulation) 10
Accumulation Steps 10
Warmup Steps 10% of update steps
$\theta_{\text{init}}$ (Bitune) 0.01
RNG Seeds 42, 43, 44
Table 10: Dataset-specific hyperparameters.
Hyperparameter UltraFeedback PIQA ARC CSQA SIQA GSM8K
Epochs 3 1 5 1 1 1
Update Steps 3000 1605 555 974 3341 747
Table 11: Learning rate for given dataset-model pair, including different PEFT variants for instruction-tuning experiments.
Model UltraFeedback PIQA ARC CSQA SIQA GSM8K
Gemma-2B 3E-4 1E-3 1E-3 1E-3 1E-3 1E-3
Gemma-7B 3E-4 3E-4 3E-4 3E-4 3E-4 3E-4
Llama2-7B 3E-4 1E-3 1E-3 1E-3 1E-3 1E-3
Llama3-8B 3E-4 1E-3 1E-3 1E-3 1E-3 1E-3
Phi-2 3E-4 1E-3 1E-3 1E-3 1E-3 1E-3
Gemma-2B (DoRA) 3E-4 - - - - -
Gemma-2B (IA3) 1E-3 - - - - -
Llama3-8B (DoRA) 3E-4 - - - - -
Llama3-8B (IA3) 1E-3 - - - - -
Table 12: Configuration of PEFT methods. All other hyperparameters have default values of HuggingFace PEFT library.
Hyperparameter Value
Rank (LoRA, DoRA) 8
Alpha (LoRA, DoRA) 1
Target Modules (All) All linear layers of MLP and Self-Attention
Feedforward Modules (IA3) All linear layers of MLP
A.3 Datasets
For all experiments we used HuggingFace Datasets (Lhoest et al., 2021) library to obtain necessary datasets.
Table 13: Table with datasets and corresponding paths, to be used with HuggingFace Datasets library.
Dataset Path
UltraFeedback openbmb/UltraFeedback
PIQA piqa
ARC-Challenge allenai/ai2_arc
CSQA tau/commonsense_qa
SIQA social_i_qa
GSM8K gsm8k
MMLU hails/mmlu_no_train
A.4 Training Speed & Memory Usage
As the method introduces two additional forward passes during training, both the training speed and the memory usage are impacted. Here we present average training times and GPU memory usage on the instruction-tuning setup with 3000 update steps (30000 actual steps, due to gradient accumulation), on a single A100 GPU, for models of two different scales - Gemma-2B & Llama3-8B. Our implementation has not been optimized, which means that e.g. training times could be improved via parallelization of two passes on the prompt. Table shows average training time, peak GPU memory usage during training, and average accuracy on downstream tasks.
Train. Time [h] Memory [GB] Acc.
Model Method
Gemma-2B LoRA 1.0 14.9 45.7
Bitune 3.1 19.8 49.7
Llama3-8B LoRA 1.7 26.6 70.4
Bitune 5.3 30.8 72.7
Using these values, one can approximate required compute to reproduce results on a given tasks, as all experiments shared the same batch size and many other hyperparameters.
A.5 Initialization of Mixing Coefficient